OpenStudy (anonymous):

Kenny goes shopping for school clothes during a tax−free weekend, and he saves eight percent sales tax. Use a percent proportion to determine how much he would save if he bought clothes worth two hundred dollars. Which of the following is true? The solution is sixteen dollars, and is a denominator in the proportion. The solution is sixteen dollars, and is a numerator in the proportion. The solution is two thousand five hundred dollars, and is a denominator in the proportion. The solution is two thousand five hundred dollars, and is a numerator in the proportion.

OpenStudy (anonymous):

use the formula of Discount: SP when D% and marked price MP is given \[Sp =\frac{ 100 -d }{100 } \times MP\] Therfore\[SP = 100 - 8 / 100 \times 200 = \frac{ 92 }{ 100 } \times 200 = $184\] He saves 200 - 184 = $16

OpenStudy (anonymous):

So what is the answer? A or B

OpenStudy (anonymous):

I guess 1.) but I don't understand what it's meant by"__menator is in proportion"..Is my guess right?

OpenStudy (anonymous):

@_lilb_to_smooth_4_U_

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

yeah :thumbs-Up: